# Self-Normalized Resets for Plasticity in Continual Learning

### v1

@@ 3.2.1 Scaled Permuted Shakespeare
We limit our experiment to 4 random seeds, scales $N\in\{1,16\}$, evaluating only SNR+L2-regularization and L2-regularization with hyperparameters $\eta=0.05$ and $\lambda=10^{-4}$, presenting our results in Table 3. We first note that as model and dataset size grow, the gap in average loss between L2-regularization and SNR+L2-regularization grows substantially. Simultaneously, we see a dramatic increase in the proportion of inactive neurons with L2-regularization. At any time step, on average $0.06\%$ of neurons are inactive at scale $N=1$ while $32.9\%$ are inactive at scale $N=16$. These results suggest that resets can play a critical role in maintaining plasticity in large-scale language models.
@@
L2 - Scale 1 $0.159\pm 0.025$ $0.152\pm 0.016$ 0.06%
SNR+L2 - Scale 1 $0.154\pm 0.008$ $0.141\pm 0.031$ 0.00%
L2 - Scale 16 $0.377\pm 0.016$ $0.410\pm 0.052$ 32.9%
SNR+L2 - Scale 16 $0.324\pm 0.028$ $0.332\pm 0.073$ $1.41\cdot 10^{-6}$%
Table 3: Average loss measured on the final epoch of each task (mean $\pm$ standard deviation over 4 seeds) for the Permuted Shakespeare problem at scales $N\in\{1,16\}$ as well as the average proportion of neurons that have not fired within the last 5 mini-batches.
@@ 4 Theoretical Analysis of Learning a Single ReLU
We analyze the average expected regret for learning a single target ReLU in order to gain intuition as to why SNR achieves lower average expected regret relative to the regularization based algorithms in our set of continual learning experiments in Section 3. At a high level, our analysis provides the following conclusions.
  1. 1.
    Learning a target ReLU with gradient descent and (an idealized) version of SNR has $O(\frac{1}{T})$ average expected regret. See Theorem 4.2 and Theorem B.2 for the full proof.
  2. 2.
    Learning a target ReLU with gradient descent and L2 regularization or L2 Init has $\Omega(1)$ average regret with probability $\frac{1}{2}$ over initializations of the network weights. See Theorem 4.1 and Theorem B.1 for the full proof.
4.1 Preliminaries
We aim to learn a single ReLU activated neuron, parameterized by the family $\Theta=\mathbb{R}^{2}$.
$f(x,v)=\sigma(\tilde{v}x+b_{v})=\begin{cases}\tilde{v}x+b_{v}&\text{if $\tilde{v}x+b_{v}\geq 0$}\\ 0&\text{if $\tilde{v}x+b_{v}<0$}\end{cases}$
We refer to $v=(\tilde{v},b_{v})$ as the target parameters and $w=(\tilde{w},b_{w})$ as our network’s parameters.
Assumption 4.1.
We sample data $x,y\in\mathbb{R}\times\mathbb{R}_{+}$ according to the distribution $\mu$ where $x\sim\text{ Uniform}(-L,L)$, for some positive constant $L$, and $y|x=f(x,v)$.
We analyze the average expected regret with respect to the squared error.
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}_{\mu}[\frac{1}{2}(\sigma(\tilde{w_{t}}x+b_{w_{t}})-\sigma(\tilde{v}x+b_{v}))^{2}]$ (1)
For notational brevity, we define
$F(w)=\mathbb{E}_{\mu}[\frac{1}{2}(\sigma(\tilde{w}x+b_{w})-\sigma(\tilde{v}x+b_{v}))^{2}]$
For convenience of analysis, we suppose that the optimization oracle $O_{t}$ performs gradient descent, rather than SGD, and has access to the oracle
$\nabla F(w)=\mathbb{E}_{\mu}[(\sigma(\tilde{w}x+b_{w})-\sigma(\tilde{v}x+b_{v}))\mathbf{1}\{\tilde{w}x+b_{w}\geq 0\}(x,1)]$
Therefore, for a learning rate $\eta$, without any regularization or resets, $\{w_{t}\}_{t=0}^{T-1}$ satisfies
$w_{t+1}=w_{t}-\eta\nabla F(w_{t})$
4.2 Gradient Descent with L2 Init and L2 Regularization
Given a regularization strength $\lambda\geq 0$, we consider the optimization oracle $O_{t}^{\text{L2 Init}}$ that performs gradient descent on the on the L2 Init regularized objective
$\bar{F}(w)=\mathbb{E}_{\mu}[\frac{1}{2}(\sigma(\tilde{w}x+b_{w})-\sigma(\tilde{v}x+b_{v}))^{2}]+\frac{\lambda}{2}||w-w_{0}||^{2}$
Then the L2 Init gradient descent update is simply
$w_{t+1}=w_{t}-\eta\nabla F(w_{t})-\eta\lambda(w_{t}-w_{0})$ (2)
We can also consider the gradient descent update with vanilla L2 regularization
$w_{t+1}=w_{t}-\eta\nabla F(w_{t})-\eta\lambda w_{t}$ (3)
@@ 4.2 Gradient Descent with L2 Init and L2 Regularization
Suppose that $x,y$ are sampled according to Assumption 4.1 and that the target parameters $v=(\tilde{v},b_{v})$ satisfy $\tilde{v}>0$ and $b_{v}\leq 0$. Then applying gradient descent with L2 Init (2) or L2 regularization (3) with strength $\lambda\geq 0$ and learning rate $\eta>0$ satisfying
$\eta$ $\leq\frac{1}{L^{2}+1}$
$\eta\lambda$ $<1$
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}_{\mu}[\frac{1}{2}(\sigma(\tilde{w_{t}}x+b_{w_{t}})-\sigma(\tilde{v}x+b_{v}))^{2}]\geq F(0)>0$

### v2

@@ 3.2.1 Scaled Permuted Shakespeare
We limit our experiment to 4 random seeds, scales 1 and 16, evaluating only SNR+L2 and L2 with hyperparameters $\eta=0.05$ and $\lambda=10^{-4}$, presenting our results in Table 3. We first note that as model and dataset size grow, the gap in average loss between L2-regularization and SNR+L2-regularization grows substantially. Simultaneously, we see a dramatic increase in the proportion of inactive neurons with L2-regularization. At any time step, on average $0.06\%$ of neurons are inactive at scale $N=1$ while $32.9\%$ are inactive at scale $N=16$. These results suggest that resets can play a critical role in maintaining plasticity in large-scale language models. See Appendix A which shows a scale 256 experiment, but run only up to 50 tasks, showing similar merits.
@@
L2 - Scale 1 0.159    (0.025) 0.152    (0.016) 0.06%
SNR+L2 - Scale 1 0.154    (0.008) 0.141    (0.031) 0.00%
L2 - Scale 16 0.377    (0.016) 0.410    (0.052) 32.9%
SNR+L2 - Scale 16 0.324    (0.028) 0.332    (0.073) $1.41\cdot 10^{-6}$%
Table 3: Average loss on the final epoch of each task for the scaled Permuted Shakespeare experiments, with means and standard deviations reported over 4 seeds.
3.2.2 Generalization
Lee et al. (2024) recently propose separating training loss from test loss in studying plasticity; the extant literature and the present work focuses largely on measuring training loss on each task. As a complement, we briefly consider measuring test loss on a holdout set for each task. We consider three setups: first, we consider the PM problem, and measure test loss on a set of 10000 image label-pairs for each task; we annotate this setup PM-G1. Our second setup is a modification of PM-G1: in each task we add noise to the labels by randomly re-labeling a fraction of the training images. We decrease the fraction of images with noisy labels from 50% on the first task to 0 on the last; we annotate this task PM-G2. Finally in analogy to PM-G2, we consider a variant of the RC problem (RC-G1) where we decay the fraction of random labels from 50% to 0 linearly across tasks. PM-G2 and RC-G1 are analogous to the setup in Lee et al. (2024).
Detailed results of these experiments are presented in Appendix A.2; we draw the following conclusions: first, SNR displays similar relative merits to competing algorithms as in our main body of experiments on training loss. Second, through a careful study of variants of the setup in PM-G2 where we alter the number of passes through the dataset relevant to each task, we point out an important issue that requires careful attention if one is to study test loss: specifically, confounding the effects of overfitting with plasticity loss.
@@ 4 Theoretical Analysis of Learning a Single ReLU
One common hypothesis that seeks to explain plasticity loss is that the network weights obtained from minimizing loss over some task yield poor initializations for a subsequent task; these initializations effectively lead to poor local minima. This section attempts to make this effect transparent through the analysis of the simplest non-trivial task: learning a single ReLU. Vardi et al. (2021) have already established that for random initializations a ReLU can be efficiently learned. Here we ask what happens when the initialization is picked adversarially; one can think of this adversarial initialization as corresponding to optimal weights for an earlier ReLU learning task.
@@
[Informal restatement of Theorems E.2 and E.1] There exists a general class of target ReLUs, a distribution over example inputs, and a distribution over initial weights for which:
  • Gradient descent with L2 Init or L2 regularization fail to learn the target ReLU with positive probability with respect to the distribution over initial weights. Specifically the average expected MSE exceeds a positive constant as the number of training examples $T$ grows with a probability bounded from below by a positive constant.